# Marathon: A Race Through the Realm of Long Context with Large Language Models

### v1

@@ 3.6 Passage Retrieval
Table 2: The evaluation results of models on Marathon benchmark. C&R refers to Comprehension and Reasoning task; MIR refers to Multiple Information Retrieval task; TR refers to Timeline Reorder task; Com. refers to Computation task; PR refers to Passage Retrieval task; SDQA refers to Short Dependency Question Answering task; Avg. denotes the average accuracy across all tasks.
4 Experiments
4.1 Setup
Models.
In this evaluation, we selected models with varying parameter sizes, ranging from 7B to 70B, and different context window sizes, ranging from 8k to 200k. The models evaluated in this study include ChatGLM3-6B-32K (Zeng et al., 2022; Du et al., 2022), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023a), Zephyr-7B-$\beta$ (Tunstall et al., 2023), Longchat-13B-16K (Li et al., 2023a), Qwen-14B-Chat (Bai et al., 2023a), Yi-34B (01.AI, 2023), Alfred-40B-1023 (Hallström et al., 2023), StableBeluga-2-70B (Mahan et al., 2023), Tulu-2-DPO-70B (Ivison et al., 2023), ChatGPT-1106 (OpenAI, 2023a), and GPT-4-1106-preview (OpenAI, 2023b).
Methods.
In this evaluation, we first assessed the inherent ability of various models to comprehend long contexts. Then, we evaluated the current mainstream methods for handling long contexts: Compression and RAG. Specifically, for the Compression method, we assessed LongLLMLingua (Jiang et al., 2023b), while for the RAG method, we evaluated two retrieval approaches using OpenAI Embedding and Jina Embedding respectively.
Type Chatglm Mistral Zephyr Longchat Qwen Yi Alfred Beluga Tulu2
Vanilla
JSON 69.48% 77.22% 84.51% 92.29% 62.29% 38.95% 13.27% 22.48% 30.72%
JSON-like 30.52% 21.18% 6.86% 3.99% 0.72% 28.56% 81.96% 0.33% 46.47%
Plain Text 0.00% 6.60% 8.63% 3.73% 36.99% 32.48% 4.77% 71.19% 22.81%
LongLLMLingua Compression
JSON 94.58% 68.10% 63.20% 93.92% 90.92% 48.10% 13.66% 29.97% 35.45%
JSON-like 5.42% 22.68% 9.15% 2.81% 0.06% 26.60% 85.95% 0.59% 43.46%
Plain Text 0.00% 9.22% 27.65% 32.68% 9.02% 25.29% 0.39% 69.54% 19.08%
OpenAI Embedding Retrieval
JSON 52.88% 84.31% 21.83% 31.11% 65.62% 67.71% 16.27% 6.27% 98.43%
JSON-like 42.42% 6.67% 69.87% 23.73% 16.93% 32.16% 83.73% 0.00% 1.11%
Plain Text 4.71% 9.02% 8.30% 45.16% 17.45% 0.13% 0.00% 93.73% 0.46%
Jina Embedding Retrieval
JSON 86.34% 83.14% 17.91% 32.75% 63.33% 65.69% 0.00% 8.43% 97.19%
JSON-like 9.15% 6.21% 73.86% 21.70% 17.91% 34.18% 100.00% 0.007% 2.16%
Plain Text 4.51% 10.65% 8.24% 45.56% 18.76% 0.13% 0.00% 91.50% 0.65%
4.2 Implementation Details
Prompt.
We used the same prompt template to ask questions for all models, and required the answers to be returned in JSON format. The specific prompt format can be seen in Figure 9.
LongLLMLingua.
For LongLLMLingua, we set the compression rate to 0.5, the dynamic context compression ratio to 0.4, We also sort the compressed contexts based on their importance.
Embedding Retreival.
For Embedding Retrieval, we utilize the ServiceContext and VectorStoreIndex of the Llama-Index (Liu, 2022). We employ various models as LLMs (Language Models), testing the OpenAI Embedding model and the Jina Embedding model as Embedding Models respectively. The default parameter settings are retained, with a chunk size of 1024 and a top-k value of 2. As for Jina Embedding, we set the pooling method to "mean" to align with Jina’s encode implementation.
Hardware.
All experiments in this evaluation were conducted on a server with 4*A100 80GB.
4.3 Results
4.3.1 Main results
Table 2 presents the performance of different models with different optimization methods on various tasks in the evaluation. For a more intuitive comparison of the results, one can refer to Figure 3, Figure 4, Figure 5, Figure 6, Figure 7, and Figure 8 in appendix. It can be observed that both OpenAI Embedding Retrieval and Jina Embedding Retrieval outperform LongLLMLingua.
In addition, all models have lower accuracy in the Timeline Reorder task and Computation task compared to other tasks. The adoption of the LongLLMLingua method did not lead to improvement, and the improvement achieved by the optimization method RAG was also limited.
Vanilla.
Among the Vanilla methods, the highest accuracy is achieved by Yi-34B, which has 34B parameters but achieves an accuracy of 55.91%. Followed by ChatGLM3-6B-32K, which achieves an accuracy of 55.05% despite having only 6B parameters. Next is Beluga-70B, which achieves an accuracy of 49.51% despite having a context window of only 4K. The average accuracy of the remaining models does not differ significantly, and none of them exceeds 40%.
LongLLMLingua.
Compared to the Vanilla method, LongLLMLingua only improved the accuracy of Qwen by 4.85%, Alfred by 1.51%, Beluga by 3.08% and Tulu2 by 8.64%. However, it reduced the accuracy of ChatGLM3 by 7.14%, Mistral by 2.8%, Zephyr by 7.74%, Longchat by 0.26%, and Yi by 7.25%.
OpenAI Embedding Retrieval.
Compared to the Vanilla method, OpenAI Embedding Retrieval improves the accuracy of Mistral by 10.37%, Zephyr by 11.66%, Qwen by 14.19%, Yi by 7.65%, Alfred by 14.05% and Tulu2 by 24.05%. However, it results in a decrease in the accuracy of ChatGLM3 by 4.06%, Longchat by 5.92% and Beluga by 1.27%.
Jina Embedding Retrieval.
Compared to the Vanilla method, Jina Embedding Retrieval has improved the accuracy of most models. Specifically, Mistral’s accuracy has improved by 12.23%, Zephyr’s accuracy has improved by 15.82%, Longchat’s accuracy has improved by 1.91%, Qwen’s accuracy has improved by 18.15%, Yi’s accuracy has improved by 7.9%, Alfred’s accuracy has improved by 13.93%, Beluga’s accuracy has improved by 6.21%, and Tulu2’s accuracy has improved by 23.60%.
4.3.2 Instruction Following Capability
During the evaluation process, we found that many models had poor instruction-following capabilities. We specified in the prompt that the large language model should respond in JSON format and provided an example of the desired response. However, sometimes the model still responded in other formats or attempted to respond in JSON format but with incomplete results. We conducted a statistical analysis of the evaluation results of various models to estimate their instruction following capabilities, the results are displayed in table 3.
JSON represents the correct generation of a JSON format response by the large language model. JSON-like refers to the attempt by the large language model to generate a response in JSON format but, for various reasons (such as exceeding length limits and being truncated, or having formatting errors), it fails to generate a correct JSON format response. Plain Text indicates that the large language model uses a different format for the response.
From Table 3, we can see that even Yi, which achieved the highest answer accuracy rate, only had an instruction-following capability of 38.95%. Beluga, on the other hand, had an instruction-following capability of only 22.48%. However, Longchat, despite having a low question-answering accuracy rate, exhibited strong instruction-following capabilities with a score of 92.29%, second only to ChatGPT’s 99.61%. Mistral and Zephyr also showed instruction-following capabilities of 77.22% and 84.51%, respectively.
5 Discussion
JSON Format.
At the recent OpenAI Developer Day [^3] https://devday.openai.com, OpenAI announced two major updates for GPT-4. One is parallel function invocation and the other is the ability to specify response format, requiring GPT-4 to return responses in JSON format. These two features are fundamental for building AGI (Artificial General Intelligence). With parallel function invocation, large language models can simultaneously call multiple utility functions to complete user tasks. The ability to respond in JSON format ensures the smooth inclusion of parameters and result returns for function invocation.
Document Context.
After the OpenAI Developer Day, GPT-4 acquired a new feature called Knowledge Retrieval, which allows users to upload their own files and retrieve relevant information from these files to respond to user queries, which is an important application scenario for RAG.
In the future, more and more large language models will implement similar functionalities. The evaluation format for long context question answering tasks will also change. Instead of directly inputting the long context in the prompt, the benchmarking work for long-context question-answering should be aligned with the direction of answering user questions based on the files provided by the user.
Online Benchmarks.
With the continuous development of large language models, the methods and benchmarks for evaluating them also need to be changed accordingly. In the past, when evaluating large language models, we relied on a research team or individual to propose a static benchmark, and then conducted evaluations based on this benchmark. However, the data in static benchmarks would leak over time and be used for model training, resulting in less accurate evaluation results. Moreover, it is inefficient for a research team or individual to build benchmarks, and it is also difficult for researchers to continuously find new evaluation data.
The power of an individual is limited, but the power of a collective is infinite. We call for the construction of Online Benchmarks, which will gather the strength of the entire open-source community to collectively maintain a continuously updated benchmark. This benchmark will have regularly updated evaluation tasks and periodically replaced evaluation formats to cope with the rapid development of large language models.
6 Conclusion
In this experiment, we compared six open source large language models with different parameter sizes and different context windows, as well as OpenAI’s ChatGPT and GPT-4. We also compared the optimization methods based on Compression, LongLLMLingua, and the optimization methods based on RAG, OpenAI Embedding Retrieval, and Jina Embedding Retrieval.
The experimental results show that the method based on RAG is more effective than the method based on Prompt Compression, and Jina Embedding performs better than OpenAI Embedding. Furthermore, the instruction-following ability of open-source large language models is generally poor, despite their ability to achieve high accuracy in question-answering tasks.

### v2

@@ 3.6 Passage Retrieval
Figure 4: An example of a test prompt. The context is truncated for display purposes.
4 Experiments
4.1 Experimental Setup
4.1.1 Models
In this analysis, we incorporated a diverse array of models, distinguished by their parameter sizes, which span from 7B to 70B, and their context window capacities, extending from 8K to 200K. Additionally, the evaluation encompassed models constructed on a state-space architectural framework. The models scrutinized in this investigation comprises ChatGLM3-6B-32K (Zeng et al., 2022; Du et al., 2022), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023a), Zephyr-7B-$\beta$ (Tunstall et al., 2023), StripedHyena-Nous-7B (Poli et al., 2023), Longchat-13B-16K (Li et al., 2023a), Qwen-14B-Chat (Bai et al., 2023a), Yi-34B (AI et al., 2024), Alfred-40B-1023 (Hallström et al., 2023), StableBeluga-2-70B (Mahan et al., 2023), Tulu-2-DPO-70B (Ivison et al., 2023), ChatGPT-1106  [^6] https://chat.openai.com, and GPT-4-1106-preview (OpenAI et al., 2024).
4.1.2 Prompt Template
Figure 4 illustrates the prompt used in our model evaluation process. The system prompt is denoted in green, the provided long context in cyan, the question related to the long context in yellow, the quartet of options in blue, and the orange segment delineates the response format for the model, accompanied by a concrete example. The instructions and response templates may vary across different models. To ensure consistency, we adjust the prompts during our evaluations to match the templates used during the models’ training phases.
4.1.3 Optimization Methods
In this evaluation, we first assessed the inherent ability of various models to comprehend long contexts. Then, we evaluated the current mainstream methods for handling long contexts: Compression and RAG. Specifically, for the compression method, we assessed LongLLMLingua (Jiang et al., 2023b), while for the RAG method, we evaluated two retrieval approaches, one based on OpenAI Embedding and the other on Jina Embedding (Günther et al., 2023).
4.1.4 Evaluation Metric
To assess the long-context comprehension abilities of large language models, we employed a multiple-choice format for the questions, each designed with several options, among which only one is the correct answer.
Accuracy=No. of Correct AnswersNo. of Total $\text{Accuracy}=\frac{\text{No. of Correct Answers}}{\text{No. of Total Answers}}$
4.1.5 Post Processing
As illustrated in the prompt template of Figure 4, we require the model to output responses in JSON format. However, current large language models are not capable of following user instructions 100% of the time. Therefore, after the model generates a response, we first parse the output using JSON format. For outputs that do not comply with the JSON format, we use regular expressions to extract the model’s responses.
4.2 Implementation Details
4.3 Vanilla
For models evaluated using the Vanilla method, we adopt the same strategy as Longbench, controlling the length of the input sequence within the context window range of the model to be evaluated by deleting the content in the middle.
4.3.1 LongLLMLingua
For LongLLMLingua, we set the compression rate to 0.5, the dynamic context compression ratio to 0.4, the condition in question to "after", and the condition compare to True. We also sort the compressed contexts based on their importance.
4.3.2 Embedding RAG
For Embedding RAG, we utilize the ServiceContext and VectorStoreIndex of the Llama-Index  [^7] https://github.com/jerryjliu/llama_index. We employ various models as LLMs (Language Models), testing the OpenAI Embedding model and the Jina Embedding model as Embedding Models, respectively. The default parameter settings are retained, with a chunk size of 1024 and a top-k value of 2. As for Jina Embedding, we set the pooling method to "mean" to align with Jina’s encode implementation.
4.3.3 Hardware
All experiments in this evaluation were conducted on a server with 4 * A100 80GB GPUs. We set the batch size to 1. For models with a scale of 7B and 13B, we use one GPU; for models around 30B, we use two GPUs; and for models at the 70B scale, we use four GPUs.
Figure 5: The instruction following the capability of different models. The x-axis represents the model, and the y-axis represents the instruction following capability. The different colors represent different methods of optimization.
4.4 Results
4.4.1 Main results
The overall accuracy of various models on the Marathon benchmark is depicted in Figure 1. Detailed performance metrics of these models, utilizing distinct optimization techniques across a range of tasks, are presented in Table 4 within the appendix. To facilitate a more comprehensive comparative analysis of the outcomes, Figures 6, 7, 8, 9, 10, and 11 are provided in the appendix. The analysis indicates that the OpenAI Embedding Retrieval and Jina Embedding Retrieval models exhibit superior performance relative to the LongLLMLingua compression.
Moreover, all examined models exhibit diminished accuracy on both the Timeline Reorder and Computation tasks relative to their performance on alternative tasks. The implementation of the LongLLMLingua failed to yield any notable enhancements, and the advancements facilitated by the RAG were similarly constrained.
4.4.2 Vanilla
Within the subset of the Vanilla method, the Yi-34B model, characterized by its 34 billion parameters, attains the highest accuracy, registering at 55.91%. This is closely followed by the ChatGLM3-6B-32K, which, despite its more modest parameter count of 6B, achieves an accuracy of 55.05%. Subsequently, the Beluga-70B model, notable for its context window limitation of 4K tokens, records an accuracy of 49.51%. The average accuracy observed across the remaining models shows minimal variance, with none exceeding the 40% threshold.
4.4.3 LongLLMLingua
In contrast to the Vanilla approach, the implementation of LongLLMLingua yielded marginal improvements in accuracy for certain models: Qwen witnessed an enhancement of 4.85%, Alfred experienced a 1.51% increase, Beluga saw a 3.08% uplift, and Tulu2 benefited from an 8.64% augmentation. Conversely, this methodology had a detrimental effect on the performance of other models: ChatGLM3 encountered a 7.14% decrement in accuracy, Mistral suffered a 2.8% reduction, Zephyr experienced a significant 7.74% decrease, StripedHyena and Longchat showed a marginal decline of 0.10% and 0.26% respectively, and Yi’s accuracy diminished by 7.25%.
4.4.4 OpenAI Embedding RAG
When juxtaposed with the baseline Vanilla methodology, the incorporation of OpenAI Embedding Retrieval notably enhances accuracy for several models: Mistral’s accuracy improved by 10.37%, Zephyr’s by 11.66%, StripedHyena’s by 16.26%, Qwen’s by 14.19%, Yi’s by 7.65%, Alfred’s by 14.05%, and Tulu2’s by an impressive 24.05%. Conversely, this approach has been observed to negatively impact the accuracy of certain models, with ChatGLM3 experiencing a 4.06% reduction, Longchat a 5.92% decrease, and Beluga a slight decline of 1.27%.
4.4.5 Jina Embedding RAG
Relative to the foundational Vanilla approach, the adoption of Jina Embedding Retrieval has led to accuracy enhancements across a majority of the evaluated models. Notably, Mistral’s accuracy experienced a 12.23% increase, Zephyr’s accuracy rose by 15.82%, StripedHyena’s accuracy increased 17.37%, Longchat saw a 1.91% improvement, Qwen’s accuracy was augmented by 18.15%, Yi’s accuracy escalated by 7.9%, Alfred’s accuracy advanced by 13.93%, Beluga’s accuracy grew by 6.21%, and Tulu2’s accuracy surged by 23.60%.
4.5 Instruction Following Capability
In our evaluation, numerous models exhibited limited ability to follow instructions accurately. We explicitly requested responses in JSON format, as exemplified by a sample provided. Nonetheless, models occasionally responded in alternate formats or attempted JSON responses that were either incomplete or incorrect. Our statistical analysis, summarized in Table 3, categorizes responses as "JSON" for correct JSON format, "JSON-like" for flawed attempts at JSON due to errors like truncation or formatting issues, and "Plain Text" for responses in other formats. For a clearer comparison of models’ ability to follow instructions after applying various optimizations, we focused on the rate of correct JSON responses as a measure of this capability, as depicted in Figure 5.
While Yi exhibited high accuracy in question answering, its compliance with instructions was notably lower, at 38.95%. In contrast, Beluga’s adherence rate to instructions was even lower, at 22.48%, despite its capabilities. On the other hand, Longchat, despite its modest accuracy in answering questions, showcased a remarkable proficiency in following instructions, achieving a 92.29% compliance rate, closely trailing behind ChatGPT’s 99.61%. The three distinct optimization techniques in our assessment demonstrated efficacy in diminishing the context length. However, it is noteworthy that none of these strategies consistently enhance the models’ ability to follow instructions.
5 Discussion
5.1 Tendency of Long Responses
During our analysis, we observed that open-source large language models often generate lengthy responses, even with clear instructions for concise JSON-formatted answers. This tendency results in the generation of extraneous content, necessitating post-processing to isolate the needed information. Table 3 presents statistics on the models’ output formats, highlighting their instruction-following capabilities. This issue likely stems from the models’ training on predominantly long responses, making it challenging for them to comply with requests for brevity.
5.2 State Space Models
Recent studies, such as Mamba (Gu and Dao, 2023), highlight the advantages of state space models (SSMs) for long context reasoning tasks. StripedHyena (Poli et al., 2023) innovatively merges SSMs with transformer structures, indicating a new direction in large language models. Despite these advancements, our analysis reveals that StripedHyena underperforms in detailed long context question answering compared to traditional transformers and does not reduce memory usage effectively, even with advanced attention mechanisms like Flash Attention 2 Dao (2023). These findings suggest the need for further optimization in State Space Models.
5.3 JSON Format
During the recent OpenAI Developer Day [^8] https://devday.openai.com, significant advancements in the capabilities of GPT-4 (OpenAI et al., 2024) were unveiled by OpenAI, notably the introduction of parallel function invocation and the specification of response formats in JSON. The parallel function invocation allows for the concurrent execution of multiple utility functions by large language models, thereby facilitating the efficient completion of complex user tasks. Moreover, the integration of JSON format for responses is instrumental in ensuring the seamless transmission of parameters and retrieval of results during function invocation, which is critical for the interoperability and functionality of AGI systems.
6 Future Work
6.1 Document as Context
Following the enhancements introduced at OpenAI Developer Day, GPT-4 (OpenAI et al., 2024) has been equipped with a Knowledge Retrieval feature. This allows the model to utilize user-uploaded documents for answering queries, marking a significant development in Retrieval-Augmented Generation (RAG) applications. This trend suggests that future large language models will likely adopt similar functionalities, impacting the evaluation methodologies for long-context question answering. Instead of embedding lengthy contexts into prompts, future benchmarks should focus on the models’ ability to extract and utilize information from user-provided documents to respond to queries. This approach necessitates a reevaluation of current benchmarks to align with these emerging capabilities.
6.2 Multi-modal Long Context
Models such as GPT4V (OpenAI et al., 2024) and Gemini (Team, 2023a) have exhibited robust capabilities in facilitating interactions that span both visual and linguistic modalities. Likewise, open-source counterparts, including LLaVA (Liu et al., 2023a) and MiniGPT-4 (Zhu et al., 2023), have demonstrated commendable performance in assessments tailored to multimodal contexts. The utility of such models extends to various real-world applications that necessitate the processing of multimodal, extensive contexts, exemplified by the comprehensive analysis and synthesis of corporate annual reports. These applications demand not only the capacity of large language models to comprehend and infer within long textual contexts but also necessitate the integration of visual understanding abilities. Presently, the open-source community is lack of benchmarks specifically designed to evaluate the proficiency of models in handling extended, multimodal contexts. Therefore, establishing a comprehensive benchmark for multimodal, long-context capabilities is of significant importance.
7 Conclusion
In this paper, we compared 10 open-source large language models, including variations in their parameter sizes and context windows, along with OpenAI’s ChatGPT and GPT-4. We assessed two prevalent optimization techniques, such as LongLLMLingua and RAG. The experimental results indicate that RAG-based optimization enhances the performance of large language models within long-context scenarios for QA-type tasks. However, the improvement is limited for tasks involving Timeline Reorder and Computation. Despite high accuracy in question-answering, these models show limited ability to follow instructions.
Acknowledgments
Min Yang was supported by National Key Research and Development Program of China (2022YFF0902100), National Natural Science Foundation of China (62376262), the Natural Science Foundation of Guangdong Province of China (2024A1515030166), Shenzhen Science and Technology Innovation Program (KQTD20190929172835662), Shenzhen Basic Research Foundation (JCYJ20210324115614039).
Limitations
Context Length Distribution.
As depicted in Figure 3, the distribution of context lengths within the Marathon benchmark exhibits a lack of uniformity. The test instances corresponding to the tasks of Comprehension and Reasoning, Multiple Information Retrieval, Computation, Short Dependency QA, and Timeline Reorder predominantly feature context lengths that are concentrated at, or below, 130K characters. Conversely, test instances with context lengths surpassing 200K characters are notably scarce.
The test instances for the Passage Retrieval task derive from the LongBench (Bai et al., 2023b) dataset, which accounts for the markedly shorter context lengths in comparison to those associated with the remaining five tasks. This discrepancy underlies the superior performance metrics achieved by all models on the Passage Retrieval task. It is our intention to revise the test instances for Passage Retrieval to ensure consistency in context lengths with the other tasks. Furthermore, our ongoing efforts are directed towards augmenting the test instances for the remaining tasks, with the objective of achieving a uniform distribution of context lengths ranging from 60K to 260K characters across all tasks.
Evaluation.
This paper presents a preliminary evaluation of optimization techniques for long contexts, which is not all-encompassing. In terms of optimization strategies, our evaluation of the Retrieval-Augmented Generation (RAG) method was limited to the employment of the OpenAI and Jina Embedding systems, exemplifying leading commercial and open-source embedding models, respectively. However, constraints related to time and financial resources precluded the examination of several advanced embedding systems, such as Voyage (Voyage.AI, 2023), Cohere (Cohere.Team, 2023), and BGE Embeddings (Xiao et al., 2023b). In the case of the Prompt Compression approach, aside from LongLLMLingua, there are other techniques like MemWalker (Chen et al., 2023a) that merit future exploration to fully assess the advantages and drawbacks of each embedding model and optimization method.
Moreover, in scenarios involving long context, while model accuracy and adherence are crucial, the speed of inference and memory demand are also vital factors to consider. This area features a variety of sophisticated optimization methods, including H2O (Zhang et al., 2023) and StreamingLLM Xiao et al. (2023a). Subsequent research will focus on evaluating the performance of these inference optimization methods in scenarios with extensive textual content, with an emphasis on their speed of inference, memory consumption, QA precision, and instruction following capability.
Ethical Considerations
Data Source and Use.
The benchmark leverages datasets that are publicly available and designated for research purposes. We have ensured that the use of these datasets adheres to their respective licenses and terms of use, emphasizing that our utilization is strictly confined to academic and research contexts.
Content Sensitivity and Bias.
Our benchmark has been meticulously curated to exclude any content that could be deemed sensitive, such as violence, discriminatory language, or adult material.
Transparency and Reproducibility.
In the spirit of fostering an open and fair research community, we will make the questions, contexts, and options of our benchmark’s test cases publicly available. However, to maintain the integrity of the evaluation process, the correct answers to the test cases will not be disclosed. Instead, we will provide an online evaluation platform where researchers can submit their models’ responses for assessment. This system is designed to ensure fairness and objectivity in the benchmarking process, allowing for an equitable comparison of different models’ capabilities.
Appendix A Dataset Construction
A.1 Data Collection
The data in the Marathon benchmark primarily originates from LongBench (Bai et al., 2023b) and LooGLE (Li et al., 2023b). Table 2 provides specific details regarding the sources of the data and the licensing information for the benchmark.
Original Dataset License C&R MIR TR Computation PR SDQA
LongBench (Bai et al., 2023b) MIT
LooGLE (Li et al., 2023b) MIT
Table 2: Sources of data for six tasks in Marathon benchmark.
A.2 Data Processing
For Comprehension and Reasoning, Multiple Information Retrieval, Short Dependency QA, Timeline Reorder and Computation tasks, we use data from LooGLE (Li et al., 2023b) to construct the test samples. For each question, we divide its corresponding long context into text segments of 12,000 tokens each and randomly select one segment to generate a distractor option, each segment is used no more than once. We send the selected long context segment $C_{i}$, the question $Q$, and the correct answer $A$ in a specific format to GPT-4, requesting it to provide a distractor option $O_{i}$ based on $C_{i}$, $Q$, and $A$.
For Passage Retrieval task, we construct test samples with Python program. We sampled data from Longbench (Bai et al., 2023b) and increased the number of passages to extend the context length.
To ensure the effectiveness and accuracy of these distractor options, we manually verify the options of each test sample.
Appendix B Model Description
ChatGLM.
The latest open-source model in the ChatGLM series, ChatGLM3-6B, as outlined by Zeng et al. (2022); Du et al. (2022), maintains several outstanding features of its predecessors, such as smooth dialogue capabilities and a low deployment threshold.
Mistral.
The Mistral-7B-Instruct-v0.1 Large Language Model (LLM), referenced by Jiang et al. (2023a), is an instructionally fine-tuned variant of the Mistral-7B-v0.1 generative text model. It has been enhanced using a diverse range of publicly accessible conversational datasets.
Zephyr.
The Zephyr-7B-$\beta$ (Tunstall et al., 2023) represents the second installment in the Zephyr series. It is a refined version of the Mistral-7B-v0.1 (Jiang et al., 2023a), specifically enhanced through training on a combination of publicly available and synthetic datasets utilizing Direct Preference Optimization (DPO).
StripedHyena.
StripedHyena (Poli et al., 2023) is the inaugural alternative model that rivals the top open-source Transformers of comparable size in both short and long-context evaluations.
Longchat.
Longchat-13b-16k (Li et al., 2023a) is an open-source chatbot developed through fine-tuning the llama-13b model Touvron et al. (2023a). It utilizes conversations shared by users on ShareGPT, applying the condensing rotary embedding technique as discussed in the blog [^9] https://lmsys.org/blog/2023-06-29-longchat/.
Qwen.
Qwen-14B-chat (Bai et al., 2023a) represents the 14-billion parameter iteration of the Qwen large language model series, abbreviated as Tongyi Qianwen, developed by Alibaba Cloud.
Yi.
The Yi-34B (AI et al., 2024) model represents the latest generation of open-source large language models, independently trained from scratch by 01.AI.
Alfred.
Alfred-40B-1023 (Hallström et al., 2023) is an enhanced version of the Falcon-40B model, featuring an expanded context length capacity of 8192 tokens.
Beluga.
StableBeluga-2-70B (Mahan et al., 2023) is a variant of the Llama2 70B model Touvron et al. (2023a), specifically fine-tuned using a dataset styled after Orca.
Tulu2.
Tulu2-DPO-70B (Ivison et al., 2023) is a refined iteration of the Llama2 model Touvron et al. (2023a), trained using a combination of publicly available, synthetic, and human datasets through the application of Direct Preference Optimization (DPO).
Appendix C Detailed Evaluation Results
Table 4 presents the detailed performance metrics of various models, utilizing distinct optimization techniques across a range of tasks. C&R refers to Comprehension and Reasoning task; MIR refers to Multiple Information Retrieval task; TR refers to Timeline Reorder task; Com. refers to Computation task; PR refers to Passage Retrieval task; SDQA refers to Short Dependency Question Answering task; Avg. denotes the average accuracy across all tasks. To provide a more intuitive comparison of the effects of different optimization approaches on the long-context comprehension and reasoning capabilities of various models across different tasks, we also illustrated Figures 6, 7, 8, 9, 10, and 11.
Appendix D Detailed Instruction Following Capability
As shown in Figure 4, we asked the models to produce results in JSON format to assess how well they follow instructions based on their output format. Table 3 summarizes the performance of 10 open-source large language models in this regard. "JSON" means the output was exactly in JSON format. "JSON-like" refers to outputs that tried to be in JSON format but included mistakes or extra text. "Plain Text" covers outputs in other formats. Since ChatGPT and GPT-4 can always gave results in JSON format, they’re not included in Table 3.
Type ChatGLM Mistral Zephyr StripedHyena Longchat Qwen Yi Alfred Beluga Tulu2
Vanilla
JSON 69.48% 77.22% 84.51% 38.43% 92.29% 62.29% 38.95% 13.27% 22.48% 30.72%
JSON-like 30.52% 21.18% 6.86% 28.95% 3.99% 0.72% 28.56% 81.96% 0.33% 46.47%
Plain Text 0.00% 6.60% 8.63% 32.61% 3.73% 36.99% 32.48% 4.77% 71.19% 22.81%
LongLLMLingua Compression
JSON 94.58% 68.10% 63.20% 50.39% 93.92% 90.92% 48.10% 13.66% 29.97% 35.45%
JSON-like 5.42% 22.68% 9.15% 12.94% 2.81% 0.06% 26.60% 85.95% 0.59% 43.46%
Plain Text 0.00% 9.22% 27.65% 36.67% 32.68% 9.02% 25.29% 0.39% 69.54% 19.08%
OpenAI Embedding RAG
JSON 52.88% 84.31% 21.83% 29.54% 31.11% 65.62% 67.71% 16.27% 6.27% 98.43%
JSON-like 42.42% 6.67% 69.87% 52.09% 23.73% 16.93% 32.16% 83.73% 0.00% 1.11%
Plain Text 4.71% 9.02% 8.30% 18.37% 45.16% 17.45% 0.13% 0.00% 93.73% 0.46%
Jina Embedding RAG
JSON 86.34% 83.14% 17.91% 24.77% 32.75% 63.33% 65.69% 0.00% 8.43% 97.19%
JSON-like 9.15% 6.21% 73.86% 56.80% 21.70% 17.91% 34.18% 100.00% 0.007% 2.16%
Plain Text 4.51% 10.65% 8.24% 18.43% 45.56% 18.76% 0.13% 0.00% 91.50% 0.65%
Figure 6: The performance of models on comprehension and reasoning task.
Figure 7: The performance of models on multiple information retrieval task.
Figure 8: The performance of models on timeline reorder task.
Figure 9: The performance of models on computation task.
Figure 10: The performance of models on passage retrieval task.
Figure 11: The performance of models on short dependency question answering task.
@@ Appendix D Detailed Instruction Following Capability
Avg. - - 55.07% 45.98% 32.37% 26.60% 84.43% 74.13% 53.10%
Table 4: The evaluation results of models on Marathon benchmark.